OpenStudy (anonymous):

Find a rectangular equation for the plane curve defined by the parametric equations. x = t - 3, y = t2 + 5

OpenStudy (anonymous):

the first step I have is y - 5 = t2, but I'm not sure if this is correct or what follows

OpenStudy (aum):

Eliminate the parameter t. x = t - 3 t = x + 3 put this in y and you will have y as a function of x.

OpenStudy (anonymous):

y= 2(x+3) +5 is the next part?

OpenStudy (aum):

\[ y = (x+3)^2 + 5 \]

OpenStudy (anonymous):

oh my apologies, it's because the variable is in front of the number it's an exponent, I thought it was tx2 instead of t^2. Okay. y = x^2+6x + 9 + 5 y= x^2 + 6x + 14 ?

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

thanks so much for the help! for future reference, if the variable is in front of the number is it always an exponent? I'm just used to it being marked with an ^.